# In-Plane Stress Waves in Layered Media: I. Non-Hermitian Degeneracies and Modal Chirality

### v1

@@ 4 Atypical scattering behavior
Figure 14: An example of scattering around EP pairs. The same unit cell as previous figures is studied here, with $s_{2}=0.1695$, $g=-0.05$. (13(a)) and (13(b)) show the scattering phase advances for an incident longitudinal wave, while (13(c)) and (13(d)) show them for shear incidence. The sharp changes are all around $\pi$ representing a change of sign as reflection amplitudes cross zero.
@@
The application of such features may be found in single sided reflectivity and/or lasing potentials as discussed earlier. As such, the amplitude of eigenvalue pairs appears to be the dominant factor. Absent of increasing the absolute value of the gain, the interaction of EP pairs appears to also lead to higher deviations from unit amplitude for eigenvalue pairs. An example of such a study is shown in Figures 13 and 14. The effective lasing (transmission amplitude above 1) and asymmetric reflection is clearly observed. Additionally, while not shown here the jump in reflection phase as it diminishes near an EP is notable, especially as its direction is quite sensitive to $s_{2}$ (not shown here). This is equivalent to the reflection parameter encircling zero in clockwise or counter-clockwise directions. While at the EP due to the low value of amplitude, this might be a rather challenging measurement, it is possible to envision this asymmetry in phase at points slightly away from EP as the basis of an approach for determination of the wave vector. Source localization using micro-structured media has been discussed in [49], and the use of EPs for enhancing their behavior is a potentially fruitful research deferred to later work.
@@ 5 Summary and conclusions
In this work we studied the propagation of coupled in-plane longitudinal and shear stress waves in layered structures. The machinery of the transfer matrix allows for exact calculations, though closed form results are prohibitive. The method allows for including linearly visco-elastic layers and even those with mechanical gain, presumably achieved through multi-physical mechanisms. Additionally, it readily allows for identification of modal features through calculation of the eigen-polarizations. The coupled physics immediately leads to the appearance of degeneracies in the spectrum, manifested as exceptional point pairs in the band structure. Though the pairs in band structure disappear from the real frequency axis with the addition of loss, for the lossless case they lead to fascinating effects in polarization (e.g. chirality associated with the direction of propagation). Scattering parameters for finite structures can also be calculated with the same tool sets, which lends itself quite readily to potential design and optimization. Furthermore, EP pairs appear also in the spectrum of the modified scattering matrix for $\mathcal{PT}$-symmetric finite structures composed of layers with balanced gain and loss. These EP pairs lead to unusual phenomena such as broken phase symmetry, transmission amplification (above 1), single-sided reflectivity, and abrupt changes in phase behavior of scattering parameters. The exploration of the potential for use in “lasing” and wave vector measurement (source localization) is deferred to future studies.
@@ Acknowledgments
The authors acknowledge National Science Foundation support for this work through Grant No. 2219087 to University of Massachusetts, Lowell. The authors wish to thank Dr. Weidi Wang for illuminating discussions throughout the course of this work.

### v2

@@ 4 Scattering behavior in $\mathcal{PT}$-symmetric media
Figure 14: An example of scattering around EP pairs. The x-axis in all subplots represents frequency (in kHz), and the y-axis shows the phase (in radians) of the scattering coefficients. The same unit cell as previous figures is studied here, with $s_{2}=0.1695$, g=0.05g=-0.05. (13(a)) and (13(b)) show the scattering phase advances for an incident longitudinal wave, while (13(c)) and (13(d)) show them for shear incidence. The sharp changes are all around $\pi$ representing a change of sign as reflection amplitudes cross zero.
@@
The application of such features may be found in single sided reflectivity and/or lasing potentials as discussed earlier. As such, the amplitude of eigenvalue pairs appears to be the dominant factor. Absent of increasing the absolute value of the gain, the interaction of EP pairs appears to also lead to higher deviations from unit amplitude for eigenvalue pairs. An example of such a study is shown in Figures 13 and 14. The wave amplification and asymmetric reflection is clearly observed. Additionally, while not shown here the jump in reflection phase as it diminishes near an EP is notable, especially as its direction is quite sensitive to $s_{2}$ (not shown here). This is equivalent to the reflection parameter encircling zero in clockwise or counter-clockwise directions. While at the EP due to the low value of amplitude, this might be a rather challenging measurement, it is possible to envision this asymmetry in phase at points slightly away from EP as the basis of an approach for determination of the wave vector. Source localization using micro-structured media has been discussed in [50], and the use of EPs for enhancing their behavior is a potentially fruitful research deferred to later work.
@@ 5 Summary and conclusions
In this work we studied the propagation of coupled in-plane longitudinal and shear stress waves in layered linear media. The machinery of the transfer matrix allows for exact calculations, though closed form results are prohibitive. The method allows for including linearly viscoelastic layers and even those with mechanical gain, presumably achieved through multi-physical mechanisms. Additionally, it readily allows for identification of modal features through calculation of the eigen-polarizations. The coupled physics immediately leads to the appearance of degeneracies in the spectrum, manifested as exceptional point pairs in the band structure. Though the pairs in band structure disappear from the real frequency axis with the addition of loss, for the lossless case they lead to fascinating effects in polarization (e.g. chirality associated with the direction of propagation). Scattering parameters for finite structures can also be calculated with the same tool sets, which lends itself quite readily to potential design and optimization. Furthermore, EP pairs appear also in the spectrum of the modified scattering matrix for $\mathcal{PT}$-symmetric finite structures composed of layers with balanced gain and loss. These EP pairs lead to unusual phenomena such as broken phase symmetry, transmission amplification (above 1), single-sided reflectivity, and abrupt changes in phase behavior of scattering parameters [51, 52]. The exploration of the potential for use in “lasing” with transmission amplification and a rigorous analysis of mode lifetime [47] and wave vector measurement (source localization) is deferred to future studies.
@@ Acknowledgments
The authors acknowledge National Science Foundation support for this work through Grant No. 2219087 to University of Massachusetts, Lowell. The authors thank Dr. Weidi Wang for illuminating discussions throughout the course of this work.